OpenStudy (anonymous):

When finding the sum of the arithmetic series 4 + 1 + -2 + -5 + -8 + -11, you will have a step of work that contains the number -3. Use complete sentences to describe how the -3 was calculated and what it means.

OpenStudy (imstuck):

The -3 is calculated by determining what is added to the first term to get to the second term. That same number is added to the second term to get to the third term, etc. Here that is a -3. 4+(-3)=1. 1+(-3)=-2. -2+(-3)=-5. The -3 is what is known as the "d" in the arithmetic sequence formula.

OpenStudy (imstuck):

\[a _{n}=a _{1}+(n-1)d\]

OpenStudy (anonymous):

Thanks!!